# SpargeAttention: Accurate and Training-free Sparse Attention Accelerating Any Model Inference

### v6

@@ 1 Introduction
Our approach. In this work, we develop SpargeAttn, a training-free sparse attention that can be adopted universally on various tasks, including language modeling and text-to-image/video, and various sequence lengths. We propose three main techniques to improve the universality, accuracy, and efficiency. First, we propose a universal sparse mask prediction algorithm, which constructs the sparse mask by compressing each block of $Q$, $K$ to a single token. Importantly, we compress selectively based on the similarity of tokens within the block, so the algorithm can accurately predict sparse masks universally across tasks. Second, we propose a sparse online softmax algorithm at the GPU warp level, which further omits some $PV$ products by leveraging the difference between global maximum values and local maximum values in online softmax. Third, we integrate this sparse approach into the 8-bit quantized SageAttention framework for further acceleration. To the best of our knowledge, SpargeAttn is the first sparse attention method that can actually accelerate across language, image, and video models without compromising accuracy.
@@ 3.1 Sparse FlashAttention
SpargeAttn adopts the tiling strategy of FlashAttention (Dao, 2024), and skip computing the blocks that are filtered out. Consider an attention operation $S=QK^{\top}/\sqrt{d},~P=\sigma(S),~O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation. Let $N$ be the sequence length and $d$ be the dimensionality of each head; the matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$ and $P$ is $N\times N$. FlashAttention proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block sizes $b_{q}$, $b_{k}$, $b_{k}$, respectively. Then, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
@@
13:    $m_{\mathrm{local}}=\mathrm{rowmax}(S_{ij});~~m_{ij}=\mathrm{max}(m_{i,j-1},m_{\mathrm{local}})$ ;   $\widetilde{P}_{ij}=\mathrm{exp}(S_{ij}-m_{ij})$ ;   $l_{ij}=e^{m_{i,j-1}-m_{ij}}+\mathrm{rowsum}(\widetilde{P}_{ij})$ ;
@@
16:     $O_{ij}[I_{w}]=\mathrm{diag}(e^{m_{i,j-1}[I_{w}]-m_{ij}[I_{w}]})^{-1}O_{i,j-1}[I_{w}]+$ $\mathrm{Matmul}(\widetilde{P}_{ij}[I_{w}],V_{j})$ ; // Paralleled by $c_{w}$ warps.
@@ A.1 Detailed Explain and results of permutation ablation
We use five distinct prompts and pre-searched hyperparameters with $l_{1}=0.05,l_{2}=0.06$ on both CogvideoX and Mochi models. The permutation are performed separately in attention operation for $Q,K,V$ after position embedding. To retain the original order of the input sequence, an inverse permutation is performed on the output of attention; for models using visual-language joint self-attention(e.g., CogvideoX), we only permute the visual tokens. When evaluating block self-similarity, we choose a block size of $128$ for query and $64$ for key, which aligns with our kernel implementation. The precision metric(L1) is evaluated using FlashAttention2 output as ground truth.

### v7

@@ 1 Introduction
Our approach. In this work, we develop SpargeAttn, a training-free sparse attention that can be adopted universally on various tasks, including language modeling and text-to-image/video, and various sequence lengths. We propose three main techniques to improve the universality, accuracy, and efficiency. First, we propose a universal sparse mask prediction algorithm, which constructs the sparse mask by compressing each block of $Q$, $K$ to a single token. Importantly, we compress selectively based on the similarity of tokens within the block, so the algorithm can accurately predict sparse masks universally across tasks. Second, we propose a sparse online softmax algorithm at the GPU warp level, which further omits some PVPV products by leveraging the difference between global maximum values and local maximum values in online softmax. Third, we integrate this sparse approach into the 8-bit quantized SageAttention framework for further acceleration. To the best of our knowledge, SpargeAttn is the first sparse attention method that can actually accelerate across language, image, and video models without compromising accuracy.
@@ 3.1 Sparse FlashAttention
SpargeAttn adopts the tiling strategy of FlashAttention (Dao, 2024), and skip computing the blocks that are filtered out. Consider an attention operation $S=QK^{\top}/\sqrt{d},~P=\sigma(S),~O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation. Let NN be the sequence length and $d$ be the dimensionality of each head; the matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$ and $P$ is $N\times N$. FlashAttention proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block sizes $b_{q}$, $b_{k}$, $b_{k}$, respectively. Then, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
@@
13:    $m_{\mathrm{local}}=\mathrm{rowmax}(S_{ij});~m_{ij}=\mathrm{max}(m_{i,j-1},m_{\mathrm{local}})$$\widetilde{P}_{ij}=\mathrm{exp}(S_{ij}-m_{ij})$$l_{ij}=e^{m_{i,j-1}-m_{ij}}l_{i,j-1}+\mathrm{rowsum}(\widetilde{P}_{ij})$;
@@
16:     $O_{ij}[I_{w}]=\mathrm{diag}(e^{m_{i,j-1}[I_{w}]-m_{ij}[I_{w}]})O_{i,j-1}[I_{w}]+$ $\mathrm{Matmul}(\widetilde{P}_{ij}[I_{w}],V_{j})$ ; // Paralleled by $c_{w}$ warps.
@@ A.1 Detailed Explain and results of permutation ablation
We use five distinct prompts and pre-searched hyperparameters with $l_{1}=0.05,l_{2}=0.06$ on both CogvideoX and Mochi models. The permutation are performed separately in attention operation for $Q,K,V$ after position embedding. To retain the original order of the input sequence, an inverse permutation is performed on the output of attention; for models using visual-language joint self-attention(e.g., CogvideoX), we only permute the visual tokens. When evaluating block self-similarity, we choose a block size of 128128 for query and 6464 for key, which aligns with our kernel implementation. The precision metric(L1) is evaluated using FlashAttention2 output as ground truth.